40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 239 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  4960   Mon Jul 11 14:03:37 2011 steveHowTolorehow to visit your old lab

Alberto is visiting us from Australia. He brought some terrific presents. It is going to be very demanding task to wait for the rest of the 40m team

to return from Wales to taste coffee:  PNG Peaberry of Wagonga, Monsooned Malabar of Jindebah and Signature Blue Blend of Cosmorex.

Attachment 1: P1080069.JPG
P1080069.JPG
  4959   Mon Jul 11 10:10:31 2011 IshwitaConfiguration AA board

The AA board shown in attachment 1 will be used in the seismometer hardware setup. A cartoon of this setup is shown in attachment 2.

BNC connectors are required for the seismometer breakout boxes. So the four-pin LEMO connectors present in the AA board were removed and panel mount BNC connectors were soldered to it. Red and blue colored wires were used to connect the BNC connectors to the board. Red wire connects the center of the BNC connector to a point on the board and that connection leads to the third leg (+IN) of the IC U### and the blue wire connects the shield of the BNC connector to the second leg (-IN) of the IC U###.

All the connections (including BNC to the AA board and in the AA board to all the filters) were tested using a multimeter by the beeping method and it was found that channel 10 (marked as C10) had a wrong connection from the point where the red wire (+ve) was connected to the third leg (+IN) of IC U91  and channel 32 (marked as C32) had opposite connections meaning the blue wire is connected to the third leg (+IN) of IC U311 and red wire is connected to the second leg (-IN) of IC U311.

Attachment 1: P7080305.JPG
P7080305.JPG
Attachment 2: seismometers.png
seismometers.png
  4958   Fri Jul 8 20:50:49 2011 sonaliUpdateGreen LockingPower of the AUX laser increased.

The ETMY laser was operating at 1.5 A current and 197 mW power.

For the efficient frequency doubling of the AUX  laser beam at the ETMY table, a higher power is required.

Steve and I changed the current level of the laser from 1.5 A to 2.1 A in steps of 0.1 A and noted the corresponding power output . The graph is attached here.

The laser has been set to current 1.8 Amperes. At this current, the power of the output beam just near the laser output is measured to be 390 mW.

The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).

The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.

Next we are going to adjust the green alignment on the ETMY and then measure the power of the beam.

At the output end of the fibre on the PSL, a power meter has been put to dump the beam for now as well as to help with the alignment at the ETMY table.

Attachment 1: Graph3.png
Graph3.png
Attachment 2: ETMY_beam_powers.png
ETMY_beam_powers.png
  4957   Fri Jul 8 19:50:19 2011 SureshUpdateRF SystemLSC rack channel assignment

[Jamie, Suresh]

   We looked at the ADC channel assignments in the LSC model and wanted to make sure that the LSC rack wiring and the LSC model are in agreement with each other.  So the plan is to wire the rack as shown below.  I will also post this file on svn so that we can keep it updated in case there are changes.

 

1Y2_Rack_Layout.png


 

  4956   Fri Jul 8 09:53:49 2011 Nicole SummarySUSSummer Progress Report 1

A copy of my summer progress report 1 has been uploaded to ligodcc 7/711 and I have just added a copy to the TTsuspension wiki

 

PDF copy of Summer Progress Report

  4955   Thu Jul 7 15:34:44 2011 JenneSummaryLSCLSC Whitening Filters have been fit

Quote:

During the daytime either tomorrow or Friday I'll adjust the actual dewhitening filters to match the measured zpk values.

 I made a handy-dandy table showing the zpk values for each whitening filter in the wiki: New whitening filter page

Next on the whitening filter to-do list: actually put these values into the dewhitening filters in foton.

  4954   Thu Jul 7 14:27:16 2011 SureshUpdateIOOsomething wrong ? : Power incident on REFL11 and REFL55

Just tying up a loose end.  The next day Kiwamu and I checked to see what the trouble was.  We concluded that the PRM had not moved during my measurement though I had 'Misaligned' it from the medm screen.  So all the power levels measured here were with the PRM aligned.  The power level change was subsequently measured and e-logged

Quote:

The measured change in the REFL DC power with and without PRM aligned seems unacceptably small.  Something wrong ?

The difference in the power with and without PRM aligned should be more than a factor of 300.

         [difference in power] = [single bounce from PRM] / [two times of transmission through PRM ]

                                          = (1-T) / T^2 ~ 310,

where T is the transmissivity of PRM and T = 5.5% is assumed in the calculation.

Also the reflectivity of MICH is assumed to be 1 for simplicity.

Quote from #4894

We now have (with the PRM misaligned):

REFL11:  Power incident = 7.60 mW ;  DC out = 0.330 V  => efficiency = 0.87 A/W

REFL55:  Power incident = 23 mW ;  DC out = 0.850 V  => efficiency = 0.74 A/W

and with the PRM aligned::

REFL11:  DC out = 0.35 V  => 8 mW is incident

REFL55: DC out = 0.975 V  => 26 mW is incident

 

 

  4953   Thu Jul 7 13:57:28 2011 steveHowToSAFETYhow not to

I found baked allen keys on the top of the clean optics cabinet.  Somewhat heavy box that can come down in an earthquake on our heads.

NOTHING SHOULD GO ON THE TOP OF THE CABINETS OR RACKS except  small plastic boxes that storing our clean clothing.

  4952   Thu Jul 7 10:25:34 2011 steveHowToGeneralcleaning out refregerator

Please ask the owner unless  it is rotten. Do not put food into garbage can inside. Take them outside so you are not inviting ants !

  4951   Thu Jul 7 02:23:59 2011 JenneSummaryLSCLSC Whitening Filters have been fit

I have fit all of the LSC whitening filters using vectfit4.m

All the data is in my folder ..../users/jenne/LSC_WhiteningTest_29June2011/

The zpk info is saved with each plot of the fit.  The pdfs are kind of huge to stitch together (or rather my computer doesn't want to do it), so I'll just post a representative one for now.

AS55Q.png

During the daytime either tomorrow or Friday I'll adjust the actual dewhitening filters to match the measured zpk values.

  4950   Wed Jul 6 23:53:38 2011 kiwamuSummaryGeneralminutes of 40m meeting
Attachment 1: meeting.pdf
meeting.pdf meeting.pdf
  4949   Wed Jul 6 23:03:57 2011 kiwamuUpdateLSCmodified locking scripts

[Jenne / Kiwamu]

 Last night we modified the locking scripts, that were called from C1IFO_CONFIGURE.adl, to adapt them to the new "PRCL" and "SRCL" convention.

So far they work fine and quitted dumping some error messages about inexistence of these channel names.

  P.S. The locking scripts have been summarized on the 40m wiki

Quote from #4912

- Now the power and signal recycling cavity lengths are named "PRCL" and "SRCL" in stead of three letter names without "L".

We should change the locking script to accomodate these changes.

  4948   Wed Jul 6 18:15:26 2011 kiwamuUpdateSUSRe: BS oplev spectra

[Steve / Kiwamu]

Motivation:

 Since the oplevs were the ones we haven't carefully tested, so the oplevs need to be checked.

This checking is also a part of the suspension optimizations (see the minutes of the last 40m meeting).

 In this work Steve will check two things for all the oplevs :

    1. Noise level including the dark noise, electrical noise and ADC noise to just make sure that the noise are blow the signal levels below ~ 30Hz.

    2. The spectra of the signals to make sure there are no funny oscillations and unexpected structures

 

Measurement :

  To check the things listed above, we take two kinds of oplves' spectra :

     1. "dark noise" when the He-Ne beam is blocked.

     2. "signals" when the optics are damped by only OSEMs

 We did these checks on the BS oplev today (see the last entry).

All of them are fine, for example the dark noise (including electrical noise and ADC noise) are below the signal levels.

And no oscillation peak was found. Steve will go through all of the oplevs in this way.

Quote from #4947

Healthy BS oplev

  4947   Wed Jul 6 16:44:37 2011 steve, kiwamuUpdateSUSBS oplev spectra

Healthy BS oplev

Attachment 1: BS.jpg
BS.jpg
  4946   Wed Jul 6 15:32:32 2011 JamieUpdateSUSRe : More normalization of all sus controllers

So after talking to Kiwamu about it, I understand now that since the damping loops need all of this extra gain when the high-pass corner is moved up, it's more convenient to put that gain in the control filter itself, rather than having to crank the overall DC gain up to some inconveniently high value.

  4945   Wed Jul 6 11:45:20 2011 JamieUpdateSUSMore normalization of all sus controllers

Quote

I'm attaching a screenshot of some of the problems I see so far with MC3.

I tried to fix all of the problems that I could identify in this screen shot:

  • Fixed the TO_COIL output filter matrix screen to correctly point to the matrix element filter screens (all SUS)
  • Removed MCL sections from SUS_XXX_POSITION screens, except for MC2.  I also modified the _POSITION screens for the ETMs to refer to ALS instead of MCL.
  • Zeroed out all of the lockin gains in the TO_COIL matrices (MC SUS)
  • Made sure all whitening filter were ON (all SUS)
  • Made sure all cts2um calibration filters were on (all SUS)
  • Made sure all oplev servos were on (all SUS)
  4944   Wed Jul 6 10:35:35 2011 JamieUpdateSUSRe : More normalization of all sus controllers

Quote:

We found the 30 Hz high pass filters had lower gain than what they used to be at low frequcnies.

So we increased the gain of the high pass filters called '30:0.0'  by a factor of 10 to have the same gain as before.

 

I'm not convinced that this is what you want to do, or at least I wouldn't do it this way.  The "k" in the zpk filter was set such that the filter had unity gain above the high-pass cut-off frequency.  For a 30 Hz high-pass the k needs to be a factor of 10 smaller than it would be for a 3 Hz high-pass to achieve this high frequency unity gain.

As it is now these HP filters have 20 dB of gain above 30 Hz.  If the open loop transfer function needs to more gain I would have done that by adjusting the overall DC gain of the filter bank, not by increasing the gain in this one filter.  Maybe you guys have been doing it differently, though.  Or maybe I'm just completely off base.

  4943   Wed Jul 6 02:12:36 2011 kiwamuUpdateSUSRe : More normalization of all sus controllers

[Jenne / Rana/ Kiwamu]

 We found the 30 Hz high pass filters had lower gain than what they used to be at low frequcnies.

So we increased the gain of the high pass filters called '30:0.0'  by a factor of 10 to have the same gain as before.

Now all the suspension shows some kind of damping. Needs more optimizations, for example Q-adjustments for all the suspensions...

Quote from #4942

This is getting closer, but with the whitening left OFF and the cts2um filter also OFF, none of the suspensions are working correctly.

 

  4942   Tue Jul 5 21:26:51 2011 ranaUpdateSUSMore normalization of all sus controllers

This is getting closer, but with the whitening left OFF and the cts2um filter also OFF, none of the suspensions are working correctly. I'm shutting down all the watchdogs until someone gets around to setting the damping gains and filters correctly.

I'm attaching a screenshot of some of the problems I see so far with MC3.

I'm going to try to get the MC suspensions working OK for tonight so that we can use them for the PRMI locking work.

Update #1: None of the MC SUS DAQ channels are found by dataviewer....SUS debugging speed reduced by 10x.  Tue Jul 05 21:38:17 2011

Update #2: POS/PIT/YAW BIAS sliders now seem to work, but are ~1000x too weak to do anything.   Tue Jul 05 21:41:38 2011

 

Attachment 1: Screenshot-1.png
Screenshot-1.png
  4941   Tue Jul 5 18:57:10 2011 JamieUpdateSUSMore normalization of all sus controllers

Based on Rana's comment I have gone through and moved all of the corner frequencies for the high pass filters in the SUS damping controllers to 30 Hz.  I did this for all optics (MC1, MC1, MC3, BS, ITMX, ITMY, PRM, SRM, ETMX, ETMY) all degrees of freedom (POS, PIT, YAW, SIDE).

Rana also suggested I turn off all of the BounceRoll filters until we get a chance to tune those individually for all the suspensions.

Finally, I normalized the MC SUSXXX filter banks to look just like all the other suspensions.

All damping filter banks for all degrees of freedom for all single suspensions should all be the same now (modulo the differences in the BounceRoll filters, which are now turned off).

  4940   Tue Jul 5 17:38:46 2011 kiwamuUpdateABSLABSL laser frequency-locked

In this past weekend the ABSL laser was successfully frequency-locked to the PSL laser with a frequency offset of about 100 MHz.

In the current setup a mixer-based frequency discriminator is used for detection of the beat-note frequency.

 


Setup for frequency locking

 The diagram below shows the setup for the frequency locking.

RFsetup.png

(UGF)
  According to a brief check of the loop oscillation it implies that the UGF is around 40 kHz.
Of course I will measure the open-loop over a wide frequency range at some point.
The lock was quite robust and it was able to stay locked for more than an hour as far as I observed.
 
(Beat-note detection and its amplification)
 The fast RFPD, which is standing on the PSL table to detect the beat-note (see the picture on #4939), showed the signal level of -16 dBm.
Then the signal goes through the RF amplifier stage to have an appropriate power level at the mixer in the frequency detection stage.
Before the signal goes into the amp. stage I put a power splitter so that I can see a signal on a spectrum analyzer.
 
(Frequency detection)
 In the mixer-based frequency discriminator, the length difference between two coax. cable was chosen to be 1 m.
This length difference gives us a zero cross point (operation point) of ~100 MHz and linear range of ~ +/-100MHz, which can sufficiently cover the FSR of PRC: 22MHz.
 
(Control filter)
 After the signal goes through the freq. discriminator, a low noise amplifier, ITHACO 1201 was installed as a control filter.
The reason I chose ITHACO 1201 was that it has a larger output range than that of SR560.
ITHACO 1201 can go to +/- 15V, which means the ABSL laser should be able to track the frequency by about +/- 65 MHz without a thermal actuation.
Right now a single pole was put at 1 Hz for an easy lock.
The output from 1201 goes directly to the laser PZT, whose input is on the front panel of the laser controller.
 

 

Temperature setpoints

 During the work I found three temperature points, where we can observe the beat-note signal within a bandwidth of 1 GHz.
 Here is a summary of the temperature set points :
    PSL temperature = 31.71 deg
    ABSL temperature = 44.19 deg (not good),
                                       47.25 deg (chosen to be nominal set point),
                                       50. 23 deg
 The first one (44.19 deg) wasn't good because changing the temperature of the ABSL laser also changes the amplitude of the beat-note significantly.
In a worst case the signal completely disappeared when the ABSL laser was at a certain temperature close to 44.19 deg.
The other two temperature points seemed good. I eventually chose the second one (47.25 deg) for the nominal set point.
Note that the current readout has been 1.81 A on the laser controller.
  4939   Tue Jul 5 16:09:54 2011 kiwamuUpdateABSLsome photos for ABSL setup

Here I show two photos of the latest ABSL (ABSolute Length measurement) setup.

APtable.png

Figure.1 : A picture of the ABSL setup on the AP table.

  The setup has been a little bit modified from the before (#4923).

 As I said on the entry #4923, the way of sampling the ABSL laser wasn't so good because the beam, which didn't go through the faraday, was sampled.

In this latest configuration the laser is sampled after the faraday with a 90% beam splitter.

The transmitted light from the 90% BS (written in pink) is sent to the PSL table through the access tube which connects the AP and PSL table .

 

PSLtable.png

FIgure.2: A picture of the ABSL setup on the PSL table.

 The 10% sampled beam ( pink beam in the picture) eventually comes to the PSL table via the access tube (the hole on the left hand side of the picture).

Then the ABSL beam goes through a mode matching telescope, which consists of a combination of a concave and a convex lens.

The PSL laser (red line in the picture) is sampled from a point after the doubling crystal.

The beam is combined at a 50 % BS, which has been setup for several purposes( see for example #3759 and #4339 ) .

A fast response PD (~1 GHz) is used for the beat-note detection.

  4938   Tue Jul 5 13:35:56 2011 steveUpdateSUSPRM damping restored

The PRM sus damping was restored. It's side rms motion came down from 35 to 4 mV immediately.      Lab   air quality is back to normal.

 

  4937   Tue Jul 5 08:32:38 2011 steveUpdatePEMbad air quality in the lab

The fireworks of yesterday showing up in the lab. Pasadena out side air 6.6 million  cfm for 0.3 micron particles  and 1.5 million cfm for 0.5 micron size.

Attachment 1: badaq.jpg
badaq.jpg
Attachment 2: july4pem.jpg
july4pem.jpg
  4936   Mon Jul 4 14:27:35 2011 kiwamuSummaryGeneralWeekly report

Summary of the week ending July 3rd.  Number of elog entries = 44 

- SUS

   * The output TO_COIL matrix were simplified
   * Checked all the BO whitening switch => Only ITMY_UL didn't switch
   * All the DOF filters were normalized.
       => All the DOF filters are now ("3:0.0", "Cheby", "BounceRoll") 
       => The High pass should have 30Hz cut off ("30:0.0") ?
   * All the resonant peaks has been fit

- LSC

   * MICH noise budget.
               => dominated by sensing noise.
   * The sensing matrix in the PRMI configuration was measured. 
               => The demodulation phase on AS55 seemed wrong. Need a doublecheck
   * A new screen, called C1LSC_OVERVIEW.adl, was released.

   * A channel name modification: "PRC" and "SRC" => "PRCL" and "SRCL" and etc.
   * The response of the LSC whitening filters were checked. 
                 => CH26 showed different phase response.
 

- MC work

 * Power budget on the AP table was made (in a high power situation).
REFL11 = 7.4 mW
REFL55 = 22 mW
        MCREFL = 114 mW
        WFS1   = 1.24 mW
        WFS2   = 2.7 mW
 * Measurement and adjustment of RFQPD response
         Resonance frequencies of WFS1 and WFS2 were adjusted. WFS1 and WFS2 were installed on the AP table
 * Started working on MCL path. 
         => needs some more CDS jobs to correctly assign ADC channels

- CDS

 * Joe modified the automated scripts for producing model webviews

- ABSL

  * The alignment of the injection beam was done.

- Fiber experiment

 * A fiber was laid down from the ETMY table to the PSL table

- TT characterization

 * The mechanical stage for the horizontal displacement measurements is set up. 

- Configuration and other topics

      * Maglev stuff has gone to bridge lab.
      * Chris.W told us that the EPICS mutex issue can be solved by upgrading the EPICS version
      * All the PDs are stored in the east arm cabinet E4
      * Safety interlocks were connected to the ETMY laser and ABSL laser
      * Cshrc.40m was modified to make 32-bit machine happy
      * NDS2 buffer size on Mafalada had been too small and was increased somewhat such that we can still work for the SUS peak fit job
 
  4935   Sun Jul 3 21:18:06 2011 ranaUpdateComputer Scripts / ProgramsstatScreen scripts dead since Feb 4 / now revived

This CSHRC mangling on Feb 4 did more than re-arrange FB binaries.

It broke the path to MEDM for the 32-bit machines in the lab (e.g. mafalda) and stopped the MEDM snapshots from being posted onto our MEDM Status Web Page.

This is because, in addition to the paths mentioned in the above elog, the paths to the EPICS directories were also commented out. I've re-inserted them into our

.cshrc file in the 32-bit section; the statScreen CRON that Yoichi set up is now back in business.

 

* for some reason, the 'cronjob.sh' script is wiping out its own log file. It would be great if someone who understands stderr output re-direction can fix it so that the log-file from each run is retained until the next time cron runs.

  4934   Fri Jul 1 20:26:29 2011 ranaSummarySUSAll SUS Peaks have been fit

         MC1    MC2    MC3    ETMX   ETMY   ITMX   ITMY   PRM    SRM    BS     mean   std
Pitch   0.671  0.747  0.762  0.909  0.859  0.513  0.601  0.610  0.566  0.747  0.698  0.129
Yaw     0.807  0.819  0.846  0.828  0.894  0.832  0.856  0.832  0.808  0.792  0.831  0.029
Pos     0.968  0.970  0.980  1.038  0.983  0.967  0.988  0.999  0.962  0.958  0.981  0.024
Side    0.995  0.993  0.971  0.951  1.016  0.986  1.004  0.993  0.973  0.995  0.988  0.019

There is a large amount of variation in the frequencies, even though the suspensions are nominally all the same. I leave it to the suspension makers to ponder and explain.

Attachment 1: Screen_shot_2011-07-01_at_8.17.22_PM.png
Screen_shot_2011-07-01_at_8.17.22_PM.png
  4933   Fri Jul 1 20:22:24 2011 ranaUpdateSUSETMY sus controller found to be in a bad state

Actually, ETMY was the only good one. They should all have the 30 Hz High pass as the damping filter. I think these details are in the elog entry that we originally made while doing ETMY.

They should all also have a 3:30 in the XXSEN to compensate the whitening. The logic is supposed to be that FM1 is ON when the hardware whitening is ON. This is the opposite of the old logic and its why the damping filter has to be moved from 3 to 30 Hz.

  4932   Fri Jul 1 18:54:34 2011 JamieUpdateSUSETMY binary whitening switching confirmed to be fixed

After finally figuring out what was messed up with ETMY I was able to get good measurements of the binary whitening switching on ETMY to determine that it is in fact working now:

ETMY
ul : 3.2937569959  = 10.3538310999 db
ll : 3.28988426634 = 10.3436124066 db
sd : 3.34670033732 = 10.4923365497 db
lr : 3.08727050163 =  9.7914936665 db
ur : 3.27587751842 = 10.3065531117 db

  4931   Fri Jul 1 18:48:13 2011 JamieUpdateSUSETMY sus controller found to be in a bad state

I'm not sure what happened to ETMY SUS, but it was in a pretty bad state.  Bad burt restore, I would guess.

Most egregiously, the inputs to all of the coil output filters were switched off.  This is a bit insidious, since these inputs being off doesn't show up on the overview screen at all.  This explains why ETMY had not been damping for the last couple of day, and why my binary whitening switching measurements were nonsense.

I also found that ETMYs damping filter was a 30 Hz high pass, instead of the 3 Hz high pass in all the other suspension controllers.  Unfortunately a messed up burt restore can't explain that.

I normalized the ETMY controller to match all of the other controllers (ie. gave it a nice new 3 Hz high pass), adjusted gains accordingly, and now ETMY is behaving nicely.

  4930   Fri Jul 1 18:41:53 2011 JamieUpdateSUSCore optic sus damping controllers normalized

I found many of the core optic (ETMs, ITMs, BS, PRM, SRM) suspension DOF damping controllers (SUSPOS, SUSPIT, SUSYAW, SUSSIDE) to be in various states of disarray:

  • Many of the controllers did not have their "Cheby" and "BounceRoll" filters switched on.
  • Some of the controllers didn't even have the Cheby or BounceRoll filters at all, or had other different filters in their place.
  • ETMY was particularly screwy (I'll make a separate follow-up post about this)
  • A bunch of random other unused filters lying around.
  • oplev servos not on
  • etc.

I went around and tried to clean things up, by "normalizing" all of the DOF damping filter banks, ie. giving them all the same filters and clearing out unused filters, and then turning on all the appropriate filters in all core optic damping filter banks ("3:0.0", "Cheby", "BounceRoll").  I also went sure that all the outputs were properly on, and the oplev servos were on.

A couple of the optics had to have their gains adjusted to compensate for filter changes, but nothing too drastic.

Everything now looks good, and all optics are nicely damped.

I didn't touch the MC sus damping controllers, but they're in a similar state of disarray and could use a once-over as well.

 

  4929   Fri Jul 1 16:01:48 2011 JamieUpdateSUSETM binary whitening switching fixed

I have fixed the binary whitening switching for the ETMs (ETMX and ETMY).  See below for a description of what some of the issues were.

The ETMX whitening/no-whitening response (same measurements performed in my previous post on checking vertex sus whitening switching) looks as it should.  The ETMY response seems to indicate that the switching is happening, but the measurements are very noise.  I had to up the averaging significantly to get anything sensible.  There's something else going on with ETMY.  I'll follow up on that in another post.

response

ETMX
ul : 3.28258088774 = 10.3243087313 db
ll : 3.31203559803 = 10.4018999194 db
sd : 3.27932572306 = 10.3156911129 db
lr : 3.28189942386 = 10.3225053532 db
ur : 3.31351020008 = 10.4057662366 db

ETMY
ul : 2.9802607099  =  9.4850851468 db
ll : 1.46693103911 =  3.3281939600 db
sd : 2.19178266285 =  6.8159497462 db
lr : 2.2716636118  =  7.1268804285 db
ur : 3.42348315519 = 10.6893639064 db

End rack cable diagrams inconsistent with binary channel mapping

One of the big problems was that the most up-to-date end rack cable diagrams (that I can find) are inconsistent with the actual binary mapping. The diagram says that:

  • BO adapter chassis output A (ch 1-16)   --> CAB_1X4_26 --> cross-connect 1X4-B7 (carrying QPD whitening switching signals)
  • BO adapter chassis output B (ch 17-32) --> CAB_1X4_27 --> cross-connect 1X4-A6 (carrying OSEM whitening switching signals)

In fact, the binary outputs are switched, such that output A carries the OSEM signals, and output B carries the QPD whitening signals.

I SWITCHED THE CABLES AT THE BINARY OUTPUT ADAPTER CHASSIS so that:

  • BO adapter chassis output A (ch 1-16)   --> CAB_1X4_27 --> cross-connect 1X4-A6 (carrying OSEM whitening switching signals)
  • BO adapter chassis output B (ch 17-32) --> CAB_1X4_26 --> cross-connect 1X4-B7 (carrying QPD whitening switching signals)

The rest of the wiring remains the same.

I made the same transformation for ETMY as well.

  4928   Fri Jul 1 11:47:25 2011 ranaUpdateIOOWFS2 resonances and installation

What is implicit in Suresh's entry is that we decided to run the WFS with the 10 dB internal attenuation set to ON as the nominal. In the past, we have always had all the attenuation OFF for max gain. The layout of the WFS is such that we get that nasty 200 MHz oscillation due to crosstalk between the 2 MAX4106 opamps for each quadrant. The 10 dB attenuator is able to reduce the positive feedback enough to damp the oscillation.

In principle, this is still OK noise-wise. I think the thermal noise of the resonant circuit should be ~2-3 nV/rHz. Then the first opamp has a gain of 5, then the -10 dB attenuator, then another gain of 5. The noise going to the demod board is then ~10-15 nV.

The real noise issue will be the input noise of the demod board. As you may recall, the output of the AD831 mixer goes to a AD797. The AD797 is a poor choice for this application. It has low noise only at high frequencies. At 10 Hz, it has an input voltage noise of 10 nV/rHz and a current noise of 20 pA/rHz. If we wanted to use the AD797 here, at least the RC filter's resistor should be reduced to ~500 Ohms. Much better is to use an OP27 and then choose the R so as to optimize the noise.

We should also be careful to keep the filter frequency low enough so as not to rate limit the OP27. From the schematic, you can see that this circuit is also missing the 50 Ohm termination on the output. There ought to be the usual high-order LC low pass at the mixer output. The simple RC is just not good enough for this application.

As a quick fix, I recommend that when we next want to up the WFS SNR, we just replace the RC with an RLC (R = 500 Ohms, L = 22 uH, C = 1 uF).

 

Attachment 1: Screen_shot_2011-07-01_at_11.13.01_AM.png
Screen_shot_2011-07-01_at_11.13.01_AM.png
  4927   Fri Jul 1 07:01:23 2011 SureshUpdateIOOWFS2 resonances and installation

This was the WFS whose photodiode was repaced as the old one was found to be damaged. 

I retuned the resonances and the notches of all the quadrant and have attached a pdf file of my measurements.

 

Some notes:

a)  The variable inductor on WFS2Q2 quadrant may need to be changed. The ferrite code has come of the solinoid and is just held in place due to friction..  It may be easily disturbed.    So though i chose to leave it in place for now,  it will need to be replace in case the Q3 misbahaves..

b) In general the frequencies have shifted a bit when I closed the lid of tne WFS sensor head.

 

WFS1 and 2 have been installed on the AP table and are functional. I am shifting attention to the software.

 

Attachment 1: WFS2new.pdf
WFS2new.pdf WFS2new.pdf WFS2new.pdf WFS2new.pdf
  4926   Thu Jun 30 21:55:16 2011 ranaConfigurationDAQNDS2 conf change

As I recently had trouble getting all of the SUS SENSOR channels at once from NDS2, I asked J.Z. for help. He found that the number of buffers on mafalda was set to only allow a small amount of data to be requested at one time.

He's going to have to figure out a more permanent fix, but for now he's increased the data buffer size to allow somewhat larger chunks to be gotten. I have made a work around in matlab, which gets smaller chunks and then cats them together.

Its in SUS/peakFit/.

Attachment 1: Untitled.png
Untitled.png
  4925   Thu Jun 30 21:00:20 2011 kiwamuUpdateABSLfine alignment done

(Just a quick report)

The fine alignment of the ABSL laser injection was successfully done.

I was able to see the DRMI fringings at the AS camera. The ABSL beam is injected from the AS port, therefore what I saw on the camera was the reflection back from the interferometer.

 

(Things to be done)

 -  A beat-note setup on the PSL table.

 - Refinement of the mode matching. The beam spot on the AS camera is a bit bigger, so I should more tightly focus the injected beam.

Quote from #4917

 - Fine alignments will be done tomorrow

  4924   Thu Jun 30 17:24:46 2011 steveUpdateSAFETYsafety interlocks

Lightwave M126-1064-700 lasers  sn415  at east end of the Y arm and  sn201 at the AP table  are connected individually  to one each EMERGENCY LASER SHUT OFF SWITCH. 

Attachment 1: P1070949.JPG
P1070949.JPG
Attachment 2: P1070948.JPG
P1070948.JPG
  4923   Thu Jun 30 14:11:55 2011 kiwamuUpdateABSLa photo of ABSL setup on the AP table

Here is a picture of the latest ABSL setup at the east part of the AP table.

ABSLsetup.png

 

(Some notes )

 - The ABSL laser is injected from the AP port.

  - A 90 % reflection BS was installed just after the NPRO, this is for sampling a 10% of the laser to the PSL table.

    However, I've just realized that this is not a nice way because the 10 % beam doesn't  go through the Faraday. Whoops.

 - A polarzser cell at the input side of the Faraday doesn't let any beam go through it for some reasons (broken ?).

    Therefore instead of having such a bad cell, a cube PBS was installed.

 -  A room was left on the table for the AS165 RFPD (green-dashed rectangular in the picture).

Quote from #4917

 - Also a picture of the setup will be uploaded in the morning.

 

  4922   Thu Jun 30 11:40:21 2011 JamieUpdateIOORe: misc. MC work

Quote:

Jamie has started to revert the "ALTPOS" effect on the MC mirrors. So far, the screens and SLOW channels have been fixed, but the fast channels still say "ALTPOS" in the dataviewer instead of "MCL".

 The framebuilder just needed to be restarted to pull in the fixed channel names.  I restarted the framebuilder and now the channels (C1:SUS-MC2_MCL_*) are showing up properly.

  4921   Thu Jun 30 11:36:54 2011 JamieUpdateSUSRe: ITMX whitening, ETMX left free swinging

Quote:

While closing up the whitening shop for the night, I noticed that the ITMX whitening state (Whitening "On") is opposite that of all other suspensions (they all have Whitening "Off").  I don't know which way is correct, but I assume they should all be the same.  Once all the whitening and BO testing is done, we should make sure that they're all the way we want them to be.

This was certainly my fault, probably left over from early debugging of my BO switch check script.  I've turned the ITMX whitening all off, to match the other suspensions.

Quote

Also, Koji and I are leaving ETMX free swinging.  That's the way we found it, presumably from Jamie's BO testing at the end station today.  We don't know what the optic's story is, so we're leaving it the way we found it.  Jamie (or whomever left it free swinging), can you please restore it when it is okay to do so?  Thanks!

Again, this was my fault.  Sorry.  I just accidentally left this off when I finished yesterday.  Much apologies.  I've turned the ETMX watchdog back on.

  4920   Thu Jun 30 08:18:08 2011 SureshUpdateIOOThe resonances and notches on WFS1 have been tuned.

As noted before the  resonances had to be tuned to the 29.5 MHz ( or rather 29.485 MHz to match with the Wenzel) and notches to twice that frequency (58.97 MHz). 

I tuned these frequencies and remeasured the transimpedance curves .  These are in the attached pdf file. 

Some notes.

1) The variable inductances on the PCB have a ferrite core which is actually ferrite powder compacted around an iron screw.  The screw serves to provide the adjustability.  However, being iron, it seems to have rusted and so the cores are stuck.  So several of the cores splintered when I tried to adjust the frequencies.

2) The WFS1 had a finger print/smudge on the face of the PD.  I drag wiped it with methanol to get rid of it.

 

WFS1 is ready to go on the table.  I am going to work on WFS2 today.

 

Attachment 1: WFS1_tuned.pdf
WFS1_tuned.pdf WFS1_tuned.pdf WFS1_tuned.pdf WFS1_tuned.pdf
  4919   Thu Jun 30 07:42:48 2011 SureshUpdateIOOAP Table Power levels

I measured the power in various beams on the AP table to check and see if any beam is having too much power. 

I am uploading two pics one is in the "high power state" and the other is the "low power state".   High power in the MC REFL PD occurs when the MC is unlocked.  In addition the WFS also will see this  hike in power. We wish to make sure that in either state the power levels do not exceed the max power that the PDs can tolerate.

 

 

Low Power state: MC locked, PRM not aligned.                                                   High Power state: MC unlocked,  PRM aligned.

 

AS-lowP_state.pdf             AS-highP_state.pdf

  4918   Thu Jun 30 06:54:07 2011 josephbUpdateCDSModified the automated scripts for producing model webviews

Dave Barker pointed out last week that the webview of our simulink model files, generated from the installed models (i.e. in /opt/rtcds/caltech/c1/target/<system name>/simLink/) was not handling libraries properly.  Essentially the web pages generated couldn't see inside library parts.

This was caused by 2 problems.  The first being the userapps not being in the matlab path when the slwebview call was done, so it couldn't even find the libraries.  The second problem is the slwebview code by default doesn't follow libraries and links, and needs a special command to be told to do so.

I added the following lines to the webview_simlink_update.m file:

addpath('/opt/rtcds/caltech/c1/core/trunk/src/epics/simLink/lib')
for sub = {'cds','isc','isi','sus','psl'}
 for spath = {'common/models','c1/models/lib'}
   addpath(['/opt/rtcds/caltech/c1/userapps/release/' sub{1} '/' spath{1}]);
 end
end

I also changed the following:

temp = slwebview(final_files{x},'viewFile',false);

became

temp = slwebview(final_files{x},'viewFile',false,'FollowLinks','on','FollowModelReference','on');

After confirming these changes worked, I have sent a corrected version to Dave and Keith.

The webview results can be found at: https://nodus.ligo.caltech.edu:30889/FE/

 

 

  4917   Thu Jun 30 03:26:40 2011 kiwamuUpdateABSLstatus update

Status update of the absolute length (ABSL) measurement:

 - To accommodate the ABSL stuff, the AS path was relocated on the AP table.

     (In this evening Jenne was able to lock MICH with AS55, so it's working fine.)

 - On the AP table all of the necessary items, including the NPRO, a Faraday, some mirrors and etc., were in place

 - The mode matching was coarsely done. The Rayleigh range looked reasonably long.

 - Fine alignments will be done tomorrow

 - Also a picture of the setup will be uploaded in the morning.

  4916   Thu Jun 30 01:50:02 2011 JenneUpdateSUSITMX whitening, ETMX left free swinging

While closing up the whitening shop for the night, I noticed that the ITMX whitening state (Whitening "On") is opposite that of all other suspensions (they all have Whitening "Off").  I don't know which way is correct, but I assume they should all be the same.  Once all the whitening and BO testing is done, we should make sure that they're all the way we want them to be.

Also, Koji and I are leaving ETMX free swinging.  That's the way we found it, presumably from Jamie's BO testing at the end station today.  We don't know what the optic's story is, so we're leaving it the way we found it.  Jamie (or whomever left it free swinging), can you please restore it when it is okay to do so?  Thanks!

  4915   Thu Jun 30 00:58:19 2011 KojiSummaryLSCLSC whitening filter test

[Jenne, Koji]

We have tested the LSC whitening filters. In summary, they show the transfer functions mostly as expected (15Hz zerox2, 150Hz pole x2).
Only CH26 (related to the slow channel "C1:LSC-PD9_I2_WhiteGain. VAL NMS", which has PD10I label in MEDM) showed different
phase response. Could it be an anti aliasing filter bypassed???

The 32 transfer functions obtained will be fit and summarized by the ZPK parameters.


Method:

The CDS system was used in order to get the transfer functions
- For this purpose, three filter modules ("LSC-XXX_I", "LSC-XXX_Q", "LSC-XXX_DC") were added to c1lsc
in order to allow us to access to the unused ADC channels. Those filter modules have terminated outputs.
The model was built and installed. FB was restarted in order to accomodate the new channels.

- Borrow a channel from ETMY UL coil output mon. Drag the cable from the ETMY rack to the LSC analog rack.
- Use 7 BNC Ts to split the signal in to 8 SMA cables.
- Put those 8 signals into each whitening filter module.

- The excitation signal was injected to C1:SUS-ETMY_ULCOIL_EXC by AWGGUI.
- The transfer functions were measured by DTT.
- The excitation signal was filtered by the filter zpk([150;150],[15;15],1,"n")
   so that the whitened output get flat so as to ensure the S/N of the measurement.

- For the switching, we have connected the CONTEC Binary Output Test board to the BIO adapter module
   in stead of the flat cable from the BIO card. This allow us to switch the individual channels manually.

- The whitening filters of 7 channels were turned on, while the last one is left turned off.
- We believe that the transfer functions are flat and equivalent if the filters are turned off.
- Use the "off" channel as the reference and measure the transfer functions of the other channels.
- This removes the effect of the anti imaging filter at ETMY.

- Once the measurement of the 7 channels are done, switch the role of the channels and take the transfer function for the remaining one channel.

Result:

- We found the following channel assignment

  • The ADC channels and the PDs. This was known and just a confirmation. 
  • The ADC channels and the WF filter on MEDM (and name of the slow channel)

- We found that the binary IO cable at the back of the whitening filter module for ADC CH00-CH07 were not connected properly.
This was because the pins of the backplane connector were bent. We fixed the pins and the connector has been properly inserted.

- CH26 (related to the slow channel "C1:LSC-PD9_I2_WhiteGain. VAL NMS", which has PD10I label in MEDM) showed different
phase response from the others although the amplitude response is identical.

Summary of the channel assignment (THEY ARE OBSOLETE - SEPT 20, 2011)

ADC                    Whitening Filter
CH  PD                 name in medm   related slow channel name for gain
---------------------------------------------------------------------------
00  POY11I             PD1I           C1:LSC-ASPD1_I_WhiteGain. VAL NMS
01  POY11Q             PD1Q          
C1:LSC-ASPD1_Q_WhiteGain. VAL NMS
02  POX11I             PD2I           C1:LSC-SPD1_I_WhiteGain. VAL NMS
03  POX11Q             PD2Q           C1:LSC-SPD1_Q_WhiteGain. VAL NMS
04  REFL11I            PD3I           C1:LSC-POB1_I_WhiteGain. VAL NMS
05  REFL11Q            PD3Q           C1:LSC-POB1_Q_WhiteGain. VAL NMS
06  AS11I              PD4I           C1:LSC-ASPD2_I_WhiteGain. VAL NMS
07  AS11Q              PD4Q           C1:LSC-ASPD2_Q_WhiteGain. VAL NMS
08  AS55I              AS55_I         C1:LSC-ASPD1DC_WhiteGain. VAL NMS
09  AS55Q              AS55_Q         C1:LSC-SPD1DC_WhiteGain. VAL NMS
10  REFL55I            PD3_DC         C1:LSC-POB1DC_WhiteGain. VAL NMS
11  REFL55Q            PD4_DC         C1:LSC-PD4DC_WhiteGain. VAL NMS
12  POP55I             PD5_DC         C1:LSC-PD5DC_WhiteGain. VAL NMS
13  POP55Q             PD7_DC         C1:LSC-PD7DC_WhiteGain. VAL NMS
14  REFL165I           PD9_DC         C1:LSC-PD9DC_WhiteGain. VAL NMS
15  REFL165Q           PD11_DC        C1:LSC-PD11DC_WhiteGain. VAL NMS
16  NC (named XXX_I)   PD5I           C1:LSC-SPD2_I_WhiteGain. VAL NMS
17  NC (named XXX_Q)   PD5Q           C1:LSC-SPD2_Q_WhiteGain. VAL NMS
18  AS165I             PD6I           C1:LSC-SPD3_I_WhiteGain. VAL NMS
19  AS165Q             PD6Q           C1:LSC-SPD3_Q_WhiteGain. VAL NMS
20  REFL33I            PD7I           C1:LSC-POB2_I_WhiteGain. VAL NMS
21  REFL33Q            PD7Q
           C1:LSC-POB2_Q_WhiteGain. VAL NMS
22  POP22I             PD8I
           C1:LSC-ASPD3_I_WhiteGain. VAL NMS
23  POP22Q             PD8Q
           C1:LSC-ASPD3_Q_WhiteGain. VAL NMS
24  POP110I            PD9I
           C1:LSC-PD9_I1_WhiteGain. VAL NMS
25  POP110Q            PD9Q
           C1:LSC-PD9_Q1_WhiteGain. VAL NMS
26  NC (named XXX_DC)  PD10I
          C1:LSC-PD9_I2_WhiteGain. VAL NMS
27  POPDC              PD10Q
          C1:LSC-PD9_Q2_WhiteGain. VAL NMS
28  POYDC              PD11I
          C1:LSC-PD11_I_WhiteGain. VAL NMS
29  POXDC              PD11Q
          C1:LSC-PD11_Q_WhiteGain. VAL NMS
30  REFLDC             PD12I
          C1:LSC-PD12_I_WhiteGain. VAL NMS
31  ASDC               ASDC
           C1:LSC-PD12_Q_WhiteGain. VAL NMS
---------------------------------------------------------------------------

Attachment 1: chans_24_31_WeirdPhase.pdf
chans_24_31_WeirdPhase.pdf
Attachment 2: Octopus.jpg
Octopus.jpg
Attachment 3: Test_Inputs_Plugged_In.jpg
Test_Inputs_Plugged_In.jpg
Attachment 4: Contec_Tester_Board.jpg
Contec_Tester_Board.jpg
  4914   Wed Jun 29 22:50:02 2011 ranaUpdateIOOmisc. MC work

Today I wanted to investigate the MC Length path situation for obscure reasons.

Jamie has started to revert the "ALTPOS" effect on the MC mirrors. So far, the screens and SLOW channels have been fixed, but the fast channels still say "ALTPOS" in the dataviewer instead of "MCL".

I also noticed that all of our old ADCU channels for diagnosing the PSL, MC, ISS, PMC ,etc. are completely AWOL. Let's blame Joe.

I think that there are probably some ADC channels available and that we'll just have to figure out what Joe intended for this. We certainly need it if we want to diagnose our PMC, ISS, FSS, MC, etc. Kiwamu tells us that the old PSL/IOO AA chassis is being used for some of the GCV signals, so its likely that we just have to do the appropriate channel name mapping in the DAQCONFIG tool.

Forging ahead with no data, I made up some filters in the MC2-MCL filter bank so that there could be a stable crossover between the laser path. I was able to turn it on and get some suppression of the FSS-FAST control signal, but there's no way to be sure without the fast channels. We gotta get Jamie to help us out once he finished the ETM BO mess.

  4913   Wed Jun 29 22:35:06 2011 NicoleSummarySUSCompleted Quad photodiode Box Circuit Diagrams

I have finished drawing the circuit diagrams for the quad photodiode boxes. Here are copies of the circuit diagram.

There are three main operation circuits in the quad photdiode box: a summing circuit (summing the contributions from the four inputs),

a Y output circuit (taking the difference between the input sums 3+2 and 1+4), and an X output circuit (taking the difference between the

input sums 3+4 and 1+2). I will complete an mini report on my examination and conclusions of the QPD circuit for the suspension wiki tomorrow.

summingcircuit.jpgQPDYcircuit.jpgQPDX_2circuit.jpg

 

  4912   Wed Jun 29 14:43:12 2011 KojiUpdateLSCLSC model updated

LSC model has been updated and running,

- Now the power and signal recycling cavity lengths are named "PRCL" and "SRCL" in stead of three letter names without "L".

- Names for the trigger monitor were fixed. They are now "C1:LSC-DARM_TRIG_MON", etc., instead of "...NORM"

- Channel order of the DC signals for PDDC_MTRX and TRIG_MTRX were changed.

It was "TRX, TRY, REFL, AS, POP, POX, POY" but now "AS, REFL, POP, POX, POY, TRX, TRY".

We should change the locking script to accomodate these changes.

  4911   Wed Jun 29 13:56:42 2011 kiwamuUpdateGeneralminutes of 40m meeting
Attachment 1: meeting.graffle.pdf
meeting.graffle.pdf meeting.graffle.pdf
ELOG V3.1.3-